# Interpretable Pre-Trained Transformers for Heart Time-Series Data

### v1

@@ 2.4.2 Atrial Fibrillation Fine-Tuning
2.5 Cross Entropy Loss vs Mean Squared Error
@@
Conventionally, time series transformer models produce a continuous output values for each token prediction and are thus trained with a mean squared error loss function [20]. However, unlike many time series prediction tasks, we were able to leverage the highly periodic nature of the PPG and ECG signals to create models with a well-defined small vocabulary. This meant that we were able to train the model in a similar fashion to a conventional large language model, with a cross-entropy loss function. It should be noted that we did also experiment with continuous output values and mean squared error as a loss function, but these results were less able to capture long range trends and importantly lacked interpretability in the attention weights. A full investigation into this finding is beyond the scope of this work.
@@
Figure 1: Generation error (in terms of absolute error) for both the PPG-PT and ECG-PT base models for a duration of half the length of the context window, evaluated over 40 unseen subjects from the bed-based balistocardiography dataset. The examples demonstrate that whilst the models were able to accurately capture features of the context, such as pulse shape and the dicrotic notch in the case of PPG, and all sections of the P-QRS-T for the ECG, large errors can come from slight temporal misalignment. (a) Two examples of model generations for PPG-PT, with context in black, ground truth in blue and model prediction in red. Both examples provide the maximum error of the prediction, with a large error occurring in the top example, and a more typical error occurring in the bottom example. (b) The median prediction error (red solid line) and interquartile range (red shaded area) for PPG-PT across all windows. (c) Two examples of model generations for ECG-PT, with context in black, ground truth in blue and model prediction in red. Both examples provide the maximum absolute error of the prediction, with two examples of large errors due to slight misalignments in prediction. (d) The median prediction error (red solid line) and interquartile range (red shaded area) for ECG-PT across all windows.
@@ 3 Generative Model Evaluation
In training, for all points in the context window the model outputs logits to which a softmax function can be applied, giving a predicted probability distribution of the next token over all possible tokens. For generation of future tokens, the final probability distribution is taken, representing the possible token values for the next token. The multinomial function in PyTorch is then used to sample this distribution and generate a predicted next token. Each new predicted token can then be appended to the previous context, and the last 500 tokens in this new context are then used to predict the next token. This process can be repeated until a maximum number of generated tokens has been reached. Another way to examine the predictive capabilities of each model would be to predict the next token as the one with the highest probability, instead of sampling from a distribution. The problem with this second approach is that sometimes in low frequency periods, such as a trough in PPG or a period between P-QRS-T regions in the ECG, taking just the token with maximum likelihood could allow the model to become stuck in predicting the sample token value over and over again. Because of this issue, the first approach of sampling from a probability distribution was chosen to evaluate the models.
@@
For the evaluation of generative capabilities of each model, the pre-processed subset of the Bed-based ballistocardiography dataset [21] [22] was used. This dataset consists of simultaneously recorded PPG and ECG from 40 subjects, sampled at 1kHz. In the preprocessed subset, the PPG was low-pass filtered with a cut off of 10Hz and the ECG was band-pass filtered between 1 and 40Hz. As with the previous datasets, we resampled the PPG to 50Hz, and the ECG to 100Hz, before converting the signals to tokens. The 750 sample windows were extracted and split into 500 samples to form the context and 250 samples to test the prediction accuracy of the model. For each prediction point, we calculated the absolute distance between the predicted token and true token. We were then able to calculate the median and inter-quartile range of absolute prediction error against the prediction horizon for both the PPG-PT and ECG-PT models. In Figure 1(a,c) example generation waveforms and maximum absolute errors are highlighted for both the PPG-PT and ECG-PT models. Observe that the errors stem not from a failure to predict the morphology of the PPG or ECG, but from slight temporal misalignment between the prediction and the ground truth, as is common with periodic physiological signals [23]. The median absolute distance between the prediction and ground truth in PPG-PT starts at 1 token and grows to 21 tokens across the 5 second (250 samples) prediction horizon, as shown in Figure 1(b). For ECG-PT, the median error starts at 1 token and grows to 3 tokens across the 2.5 second (250 samples) prediction horizon, as shown in Figure 1(d). It should be noted that we expect the errors to be significantly higher in the PPG-PT model when compared with the ECG-PT model, as due to the comparatively broader shape of the peaks in the PPG waveform, errors in temporal misalignment increase the prediction error across the whole waveform. This is in contrast to ECG where the vast majority of the signal is flat, and thus slight temporal misalignment is mainly penalised around the QRS complex and the T wave.
@@
Whilst these prediction errors show us that the models are good at extrapolating PPG and ECG waveforms into the future, this can never be perfect given that there may be changes in heart rate, respiration, and even arrhythmia events that are impossible to predict from the context. The important result here is the ability of the models to generalise to unseen distinct morphologies of PPG and ECG as context. It is clear that the PPG-PT condenses knowledge of peak width and the shape of the dicrotic notch (the sub-peak) in the PPG examples and is thus able to generate PPG with the same properties. This is even more obvious in the case of ECG, where both examples have very different P, Q, R, S and T wave morphologies, and ECG-PT was able to replicate this well. This is the first indicator that the PPG-PT and ECG-PT models clearly pay attention to the relevent features of both signals.
4 Interpretability of aggregate model attention
@@
Figure 2: Aggregate attention across all attention heads for the attention row corresponding to the prediction point, per transformer layer shown (the first and last layers), for both the PPG-PT and ECG-PT models. The attention maps demonstrate that in order to predict the next token, the models first look at all tokens in the immediate cycle to gain an understanding of where a token falls within that cycle. When this understanding has been formed, models then look at the all similar points occuring in other cycles in the context window. For both models, attention in the first transformer layer is shown on top and the last transformer layer (bottom), with context (black line) and the prediction point (blue circle) then corresponding to overlaid transformer attention (red transparent bars), with transparency scaled based on the attention weights shown below (red solid line). (a) The PPG-PT aggregate attention, for the prediction of a peak in the given previously unseen photoplethymography context. (b) The ECG-PT aggregate attention, for the prediction of a peak in the given previously unseen electrocardiography context.
@@
The next step to interpreting the aggregate attention of the model is to examine the attention weights of different transformer blocks for specific context examples. In Figure 2 (a) we examine the summed attention weights in PPG-PT from the row associated with predicting the peak point highlighted with a blue circle. As expected, in order to predict the peak point the first transformer block focuses on the immediate points before the peak point in order to learn how the point fits in to the local PPG cycle. This also is true of the attention matrix rows in the first transformer block associated with all previous tokens in the context. Importantly, observe that in the final transformer block, in order to predict the peak point the network focuses its attention on all previous peaks in the context. This is also true of the ECG-PT model, as demonstrated in Figure 2 (b).
5 Vector similarities between points of interest
@@
Figure 3: Plots showing the cosine similarity of points in rising slopes (blue transparent circles, blue solid lines) and falling slopes (red transparent circles, red solid lines) with similar input values, upon propagation through the transformer layers of the model. In each case a single point on a rising slope is chosen as a comparison point, represented by a black cross on the context (black solid line), and a black dotted line with a cosine similarity of 1 in the cosine similarity plot. The point of propagation through the model is highlighted with grey transparent blocks labeled with the corresponding transformer layer (from 1 to 8). In both examples, rising slope points and falling slope points are shuffled with each other in embedding space upon input to the first model layer, and gradually divide into two clusters as they propagate through the model. (a) The cosine similarity in embedding space of rising slopes and falling slopes in a photoplethysmography signal, upon propagation through the PPG-PT model. (b) The cosine similarity in embedding space of rising slopes and falling slopes in the T wave of an electrocardiography signal, upon propagation through the ECG-PT model.
@@
In the previous section, we demonstrated that the pre-trained transformer models logically focus on similar points in previous cycles of PPG or ECG when trying to predict the next token. This strongly indicates that the models are able to distinguish different points in the cycle of PPG and ECG waveforms. In this section, we aim to solidify this by examining how the cosine similarity of embedded tokens with the same value, which occur at different distinct points in the PPG and ECG cycle, change upon propagation through the model. To investigate this, we chose tokens at similar points on rising slopes and falling slopes on the PPG signal, and on the T-wave of the ECG signal. The similarity of these tokens was calculated with respect to the final rising slope in each context window. If the models are effective at distinguishing different points in the cycle, we would expect rising slopes to cluster with each other in vector space, and falling slopes to form a separate cluster in vector space. The parallel of this experiment in a large language model would be to look at the vector similarity of homonyms (words that have the same spelling but can have multiple different meanings), and examine how the vector similarity changes based on specific context in a sentence. For example, we would expect the token “bat" to form separate clusters in vector space upon propagation through an LLM based on if it contextually refers to a club to hit a cricket ball or refers to the flying mammal.
@@
6 Attention Maps of Individual Attention Heads
Figure 4: Attention maps of the attention weights of individual attention heads of last layer of both the PPG-PT and ECG-PT models. Averaged attention maps are generated by averaging the final layer of the attention weight matrix over the prediction of the next 500 tokens beyond the initial context window. The model context is plotted in black, with attention weights plotted in red, pink, and blue for different attention heads. Peaks in attention are highlighted in the same colour with transparent circles, and overlaid on the context with transparent bars. (a) Mapping of the averaged attention weights in the final layer attention heads of the PPG-PT model for an example photoplethysmography context. It is clear that the 3rd attention head (red) looks primarily for peaks in PPG, the 4th attention head (pink) looks primarily for troughs in PPG, and the 6th (blue) looks for the dicrotic notch. (b) Mapping of the averaged attention weights in the final layer attention heads of the ECG-PT model for an example electrocardiography context. The 1st attention head (red) looks for R peaks in the ECG, the 5th attention head (pink) looks primarily for P waves, and the 7th attention head (blue) looks primarily for the Q portion of the QRS complex of ECG.
@@ 6 Attention Maps of Individual Attention Heads
For each model, upon inputting the context, we looked at the final row in the attention weights of each head in the last transformer block, corresponding to the weights associated with predicting the next token. Upon saving these weights, we inputted a new context corresponding to the original shifted by 1 sample into the future. We then saved these weights and added the final N-1 weights to the last N-1 original weights. We repeated this shift and add operation a further N-2 times, and divided each of the N weights element-wise by the number of additions made to them. This gave us an average map of attention to the context for the next 500 token predictions, for each self-attention head in the final transformer block. We then used the findpeaks function in MATLAB, with a minimum peak distance of 15 tokens and a minimum peak height of 0.5 of the maximum attention weight in the window, to highlight the local maximas in attention.
Figure 4 highlights that individual attention heads do indeed pay attention to important features in the signal of interest. In particular, PPG-PT has an attention head that looks for peaks, another that looks for troughs and another which looks for the dicrotic notch. In addition, we find that ECG-PT has a head which looks for the Q portion of the QRS, another which looks for R-peaks and another that looks for P waves. Therefore, on top of an ability to distinguish all separate points in a PPG and ECG cycle, the pre-trained models presented in this paper also pay specific attention to some of the most physiological relevent features.
7 Change in Attention when Screening For Atrial Fibrillation
@@
Through pre-training we have obtained models which have a strong grasp on the important features of PPG and ECG, and generalise well to unseen morphologies. However, rather like in LLMs, predicting the next token of PPG and ECG lacks utility. To fully make use of these pre-trained models, we fine tune the models to classify atrial fibrillation (AF), as outlined in Section 2.4.2. Atrial fibrillation is the most common arrhythmia, characterised by an irregular heart rate where there can be rapid increases in heart rate and periods where it slows down dramatically [24]. The fine-tuned AF-PPG-PT model achieved a leave-one-subject-out area under the curve (AUC) of 0.93 when cross validated across all test subjects, and 0.96 when two subjects with poor signal to noise ratio were excluded. The fine-tuned AF-ECG-PT model achieved an AUC of 0.99 when cross validated across all test subjects. It is worth reiterating that these results were achieved with 11 minutes of fine-tuning computer time, corresponding to roughly 0.2% of the total time it took to train the base models.
@@
The key result of fine-tuning is not just the accuracy, but the interpretability of the results. To ascertain why a model made a classification, we simply aggregated the final row of soft-maxed attention weights across all heads in the final transformer layer for the base model and compared with the fine-tuned model. Any increases in attention weights therefore indicates that the model deems the corresponding tokens valuable for classification of AF. In Figure 5(a) two representative attention maps are displayed for the classification of a PPG context as AF. It is clear in both cases that the model increases attention on periods where beats occur later than expected on earlier than expected, which is the obvious characteristic of AF. This is even clearer in the fine-tuned ECG attention, given that peaks in ECG are more precisely localised in time. In Figure 5(b), an example is highlighted where a beat is expected to occur based on the previous beat timing, but it does not and thus the model attention spikes. A second example is also highlighted, where based on the previous beat timing a beat occurs much earlier than expected, and model attention therefore spikes exactly at this point. These interpretable maps of shifts in attention, demonstrating the reason why the model has made the classification, can easily be provided along with with the probability of AF.

### v2

@@ 2.5.2 PPG Beat Detection
2.6 Cross Entropy Loss and Next Token Prediction vs Mean Squared Error
@@
Conventionally, time series transformer models produce a continuous output values for each token prediction and are thus trained with a mean squared error (MSE) loss function [20] or a combination of likelihood and MSE [21]. However, unlike many time series prediction tasks, we were able to leverage the highly periodic nature of the PPG and ECG signals to create models with a well-defined small vocabulary. This allowed us to train the model exclusively with a cross-entropy loss function, in a similar fashion to a conventional large language model. It should be noted that we did also experiment with continuous output values and mean squared error as a loss function, but these results were less able to capture long range trends and importantly lacked interpretability in the attention weights. A full investigation into this finding is beyond the scope of this work.
Through training for next token prediction in combination with cross entropy loss, for each part of the context window, we force tokens to update in a way which is still relevant to that specific token. For example, tokens around peaks update to become more “peak-like” as they propagate through the transformer blocks, and the same principle is the case for all other positions in each cardiac cycle. In that sense, the final layer of attention operates on a ECG or PPG signal where each token is rich in period-specific information, and attention can thus be fine-tuned in a way that is more easily interpretable to a clinician.
@@
Figure 1: Generation error (in terms of absolute error) for both the PPG-PT and ECG-PT base models for a duration of half the length of the context window, evaluated over 40 unseen subjects from the bed-based balistocardiography dataset. The examples demonstrate that whilst the models were able to accurately capture features of the context, such as pulse shape and the dicrotic notch in the case of PPG, and all sections of the P-QRS-T for the ECG, large errors can come from slight temporal misalignment. (a) Two examples of model generations for PPG-PT, with context in black, ground truth in blue and model prediction in red. Both examples provide the maximum error of the prediction, with a large error occurring in the first, and a more typical error occurring in the second example. (b) The median prediction error (red solid line) and interquartile range (red shaded area) for PPG-PT across all windows. (c) Two examples of model generations for ECG-PT, with context in black, ground truth in blue and model prediction in red. Both examples provide the maximum absolute error of the prediction, with two examples of large errors due to slight misalignments in prediction. (d) The median prediction error (red solid line) and interquartile range (red shaded area) for ECG-PT across all windows.
@@ 3 Generative Model Evaluation
In training, for all points in the context window the model outputs logits to which a SoftMax function can be applied, giving a predicted probability distribution of the next token over all possible tokens. For the generation of future tokens, the final probability distribution is taken, representing the possible token values for the next token. The multinomial function in PyTorch is then used to sample this distribution and generate a predicted next token. Each new predicted token can then be appended to the previous context, and the last 500 tokens in this new context are then used to predict the next token. This process can be repeated until a maximum number of generated tokens has been reached. Another way to examine the predictive capabilities of each model would be to predict the next token as the one with the highest probability, instead of sampling from a distribution. The problem with this second approach is that sometimes in low frequency periods, such as a trough in PPG or a period between P-QRS-T regions in the ECG, taking just the token with maximum likelihood could allow the model to become stuck in predicting the sample token value over and over again. Because of this issue, the first approach of sampling from a probability distribution was chosen to evaluate the models.
@@
For the evaluation of generative capabilities of each model, the pre-processed subset of the Bed-based ballistocardiography dataset [22] [23] was used. This dataset consists of simultaneously recorded PPG and ECG from 40 subjects, sampled at 1kHz. In the preprocessed subset, the PPG was low-pass filtered with a cut off of 10Hz and the ECG was band-pass filtered between 1 and 40Hz. As with the previous datasets, we resampled the PPG to 50Hz, and the ECG to 100Hz, before converting the signals to tokens. The 750-sample windows were extracted and split into 500 samples to form the context and 250 samples to test the prediction accuracy of the model. For each predicted token, we calculated the absolute distance between the predicted token and the true token. We were then able to calculate the median and inter-quartile range of absolute prediction error against the prediction horizon for both the PPG-PT and ECG-PT models. In Figure 1(a,c), example generation waveforms and maximum absolute errors are highlighted for both the PPG-PT and ECG-PT models. Observe that the errors stem not from a failure to predict the morphology of the PPG or ECG, but from slight temporal misalignment between the prediction and the ground truth, as is common with periodic physiological signals [24]. The median absolute distance between the prediction and ground truth in PPG-PT starts at 1 token and grows to 21 tokens across the 5 second (250 samples) prediction horizon, as shown in Figure 1(b). For ECG-PT, the median error starts at 1 token and grows to 3 tokens across the 2.5 second (250 samples) prediction horizon, as shown in Figure 1(d). It should be noted that we expect the errors to be significantly higher in the PPG-PT model when compared with the ECG-PT model, as due to the comparatively broader shape of the peaks in the PPG waveform, errors in temporal misalignment increase the prediction error across the whole waveform. This is in contrast to ECG where the vast majority of the signal is flat, and thus slight temporal misalignment is mainly penalised around the QRS complex and the T wave.
@@
Whilst these prediction errors show us that the introduced models are good at extrapolating PPG and ECG waveforms into the future, this can never be perfect given that there may be changes in heart rate, respiration, and even arrhythmia events that are impossible to predict from the context. The important result here is the ability of the models to generalise to unseen distinct morphologies of PPG and ECG as context. It is clear that the PPG-PT condenses knowledge of peak width and the shape of the dicrotic notch (the sub-peak) in the PPG examples and is thus able to generate PPG with the same properties. This is even more obvious in the case of ECG, where both examples have very different P, Q, R, S and T wave morphologies, and ECG-PT was able to replicate this well. This is the first indicator that the PPG-PT and ECG-PT models pay attention to the relevent features of both signals.
4 Interpretability of the Pre-Trained Models
Figure 2: Aggregate attention across all attention heads for the attention row corresponding to the prediction point, per transformer layer shown (the first and last layers), for both the PPG-PT and ECG-PT models. The attention maps demonstrate that in order to predict the next token, the models first look at all tokens in the immediate cycle to gain an understanding of where a token falls within that cycle. When this understanding has been established, models then look at the all similar points occuring in other cycles in the context window. For both models, attention in the first transformer layer is shown on top and the last transformer layer beneath, with context (black line) and the prediction point (blue circle) corresponding to overlaid transformer attention (red transparent bars), with transparency scaled based on the attention weights shown below (red solid line). (a) The PPG-PT aggregate attention, for the prediction of a peak in the given previously unseen photoplethysmography context. (b) The ECG-PT aggregate attention, for the prediction of a peak in the given previously unseen electrocardiography context.
4.1 Interpretability of aggregate model attention
@@
The next step to interpreting the aggregate attention of the model is to examine the attention weights of different transformer blocks for specific context examples. In Figure 2 (a) we examine the summed attention weights in PPG-PT from the row associated with predicting the peak point highlighted with a blue circle. As expected, in order to predict the peak point the first transformer block focuses on the immediate points before the peak point in order to learn how the point integrates to the local PPG cycle. This also is true of the attention matrix rows in the first transformer block that are associated with all previous tokens in the context. Importantly, observe that in the final transformer block, in order to predict the peak point the network focuses its attention on all previous peaks in the context. This is also true of the ECG-PT model, as demonstrated in Figure 2 (b).
4.2 Vector similarities between points of interest
@@
Figure 3: Plots showing the cosine similarity of points in rising slopes (blue circles, blue solid lines) and falling slopes (red circles, red solid lines) with similar input values, upon propagation through the transformer layers of the model. In each case, a single point on a rising slope is chosen as a comparison point, represented by a black cross on the context (black solid line), and a black dotted line with a cosine similarity of 1 in the cosine similarity plot. The point of propagation through the model is highlighted with grey transparent blocks labeled with the corresponding transformer layer (from 1 to 8). In both examples, rising slope points and falling slope points are shuffled in embedding space upon input to the first model layer, and gradually divide into two clusters as they propagate through the model. (a) The cosine similarity in embedding space of rising slopes and falling slopes in a photoplethysmography signal, upon propagation through the PPG-PT model. (b) The cosine similarity in embedding space of rising slopes and falling slopes in the T-wave of an electrocardiography signal, upon propagation through the ECG-PT model.
@@
In the previous section, we have demonstrated that the pre-trained transformer models naturally focus on similar points in previous cycles of PPG or ECG when predicting the next token. This strongly indicates that the models are able to distinguish different points in the cycle of PPG and ECG waveforms. In this section, we aim to solidify this finding by examining how the cosine similarity of embedded tokens with the same value, which occur at different distinct points in the PPG and ECG cycle, change upon propagation through the model. To this end, we chose tokens at similar points on rising slopes and falling slopes on the PPG signal, and on the T-wave of the ECG signal. The similarity of these tokens was calculated with respect to the final rising slope in each context window. If the models are effective at distinguishing different points in the cycle, we would expect rising slopes to cluster with each other in vector space, and falling slopes to form a separate cluster in vector space. An analogous experiment in a large language model would be to look at the vector similarity of homonyms (words that have the same spelling but can have multiple different meanings), and examine how the vector similarity changes based on specific context in a sentence. For example, we would expect the token “bat” to form separate clusters in vector space upon propagation through an LLM based on if it contextually refers to a club to hit a cricket ball or it refers to the flying mammal.
@@
Figure 4: Attention maps of the attention weights of individual attention heads of the last layer of both the PPG-PT and ECG-PT models. Averaged attention maps are generated by averaging the final layer of the attention weight matrix over the prediction of the next 500 tokens beyond the initial context window. The model context is plotted in black, with attention weights plotted in red, pink, and blue for different attention heads. Peaks in attention are highlighted in the same colour with circles, and overlaid on the context with bars. (a) Mapping of the averaged attention weights in the final layer attention heads of the PPG-PT model for an example photoplethysmography context. Observe that the 3rd attention head (red) looks primarily for peaks in PPG, the 4th attention head (pink) looks primarily for troughs in PPG, and the 6th (blue) looks for the dicrotic notch. (b) Mapping of the averaged attention weights in the final layer attention heads of the ECG-PT model for an example electrocardiography context. The 1st attention head (red) looks for R peaks in the ECG, the 5th attention head (pink) looks primarily for P waves, and the 7th attention head (blue) looks primarily for the Q portion of the QRS complex of ECG.
4.3 Attention Maps of Individual Attention Heads
@@
For each model, upon inputting the context, we looked at the final row in the attention weights of each head in the last transformer block, corresponding to the weights associated with predicting the next token. Upon saving these weights, we inputted a new context corresponding to the original shifted by 1 sample into the future. We then saved these weights and added the final N-1 weights to the last N-1 original weights. We repeated this shift and add operation for a further N-2 times, and divided each of the N weights element-wise by the number of additions made to them. This gave us an average map of attention to the context for the next 500 token predictions, for each self-attention head in the final transformer block. We then used the findpeaks function in MATLAB, with a minimum peak distance of 15 tokens and a minimum peak height of 0.5 of the maximum attention weight in the window, to highlight the local maximas in attention.
Figure 4 highlights that individual attention heads do indeed pay attention to important features in the signal of interest. In particular, PPG-PT has an attention head that looks for peaks, another that looks for troughs and another which looks for the dicrotic notch. In addition, we found that ECG-PT has a head which looks for the Q portion of the QRS, another which looks for R-peaks and another that looks for P waves. Therefore, on top of an ability to distinguish all separate points in a PPG and ECG cycle, the pre-trained models presented in this paper also pay specific attention to some of the most physiological relevent features.
5 Fine-Tuning Results
@@
5.1 Change in Attention when Screening For Atrial Fibrillation
@@
Through pre-training we have obtained models which have focus on the important features of PPG and ECG, and generalise well to unseen morphologies. However, rather like in LLMs, predicting the next token of PPG and ECG lacks utility. To make full use of these pre-trained models, we fine tune the models to classify atrial fibrillation (AF), as outlined in Section 2.4.2. Atrial fibrillation is the most common arrhythmia, characterised by an irregular heart rate which can manifest itself in rapid increases in heart rate and periods where it slows down dramatically [25]. The fine-tuned AF-PPG-PT model achieved a leave-one-subject-out area under the curve (AUC) of 0.93 when cross validated across all test subjects, and 0.96 when two subjects with poor signal to noise ratio were excluded. The fine-tuned AF-ECG-PT model achieved an AUC of 0.99 when cross validated across all test subjects. It is worth reiterating that these results were achieved with 11 minutes of fine-tuning computer time, corresponding to roughly 0.2% of the total time it took to train the base models.
@@
The key result of fine-tuning is not just the accuracy, but the interpretability of the results. To ascertain why a model made a classification, we simply aggregated the final row of SoftMaxed attention weights across all heads in the final transformer layer for the base model and compared with the fine-tuned model. Any increases in attention weights therefore indicate that the model deems the corresponding tokens valuable for classification of AF. In Figure 5(a) two representative attention maps are displayed for the classification of a PPG context as AF. Observe in both cases that the model increases attention on periods where beats occur later than expected on earlier than expected, which is the obvious characteristic of AF. This is even clearer in the fine-tuned ECG attention, given that peaks in ECG are more precisely localised in time. In Figure 5(b), an example is highlighted where a beat is expected to occur based on the previous beat timing, but it does not and thus the model attention spikes. A second example is also highlighted, where based on the previous beat timing a beat occurs much earlier than expected, and model attention therefore spikes exactly at this point. These interpretable maps of shifts in attention, demonstrating the reason why the model has made the classification, can easily be provided along with with the probability of AF.
5.2 Beat Detection and Signal Quality Estimation in PPG
Figure 6: Example results of fine-tuning PPG-PT for beat detection. The input context PPG signal is plotted in black, predicted beats are highlighted with red crosses and ECG derived beats (ground truth) are designated with blue circles. The beat confidence is overlaid as vertical lines, with green indicating high confidence, orange indicating medium, and red indicating low. Both an example of high signal quality (top) and an example of low signal quality (bottom) are shown.
Evaluation of the fine-tuned PPG-PT model for beat detection on the “MIMIC perform testing dataset” yielded a median F1 score of 98.04% across subjects, which is comparable to the F1 scores of qppg [26] (96.9%) and MSPTD [27] (97.5%) as found by Charlton et al. [18]. It was found during training that whilst the model was able to accurately distinguish beats in examples with high signal quality after just 50 to 100 iterations, training it for longer was necessary to detect beats in examples with visibly poor signal quality. In addition to predicting the location of beats, our fine-tuned PPG-PT model also provides a continuous value between 0 and 1 at each location. This can be thought of as a “quasi-confidence estimate” in the prediction that a beat is indeed present at that location. Given that test dataset is clinical data, and thus the majority of the data is high quality, values in the bottom 10% of confidence were labelled as low confidence, values between 10% and 25% were labelled as medium confidence, and values in the top 75% were labelled as high confidence.
Observe in Figure 6 that in both the cases of high signal quality and low signal quality, our fine-tuned beat detector was able to correctly identify all PPG beats, within 150ms of the ECG derived ground truth beats. Importantly, the confidence estimates also aligned with signal quality, with lower confidence in areas of poor signal quality. The confidence estimates also carry over to the previous AF dataset. Beat detection was applied independently to “MIMIC Perform AF” and it was found that two subjects with the lowest classification accuracy, and visibly poor PPG signal quality, also had the lowest mean beat confidence. Conversely, the subject with the highest mean classification accuracy had the highest signal quality, as determined by the beat confidence.
It should be emphasised that the PPG-PT model was pretrained on finger-based clinical data, and thus has also been fine-tuned on clinical data. Wearable sensor data, such as from the wrist [28] or the ear [29] have different morphology to that of the finger. It is likely that further training would be necessary for these models to be transferable to wearable PPG data.